Rewrite the equation 2x – 3y = –6 as a function of x

isn't it

f(x) = 2/3 + 6

Answer by checkley71(8403)

2X-3Y=-6 OR 2X=3Y-6 OR X=3/2Y-3 SO THE F(X)=3/2Y-3